Write the equation in standard form using only integers:

x/4=y/5-7

Answer by stanbon(57967)

Write the equation in standard form using only integers:
x/4=y/5-7
---
Multiply thru by 20 to get:
5x = 4y - 140
---
Standard Form:
5x-4y = -140
================
Cheers,
Stan H.
==========